OpenStudy (anonymous):

Use the function rule f(x) = -x + 5. Find the output.

OpenStudy (freckles):

the output value for each input x will be -x+5

OpenStudy (freckles):

example the output value for input value x=-2 is -x+5=-(-2)+5=2+5=7

OpenStudy (anonymous):

No the multiple choice is A- 10 B- 7 C- -10 D- 3

OpenStudy (freckles):

you never gave any input value